Use Cumulative Standardized Normal Probability table to find the value z* for which:

a) P( Z ≤ z*) = 0.95

b) P( Z ≤ z*) = 0.2

c) P( Z ≤ z*) = 0.25

d) P( Z ≥ z*) = 0.9

e) P( 0 ≤ Z ≤ z*) = 0.41

f) P( -z* ≤ Z ≤ z*) = 0.88


1
Expert's answer
2021-03-11T14:08:23-0500

"\text{Using the table of the values of Laplace function }\\
\Phi(z)=\frac{1}{\sqrt{2\pi}}\int_0^z e^{-\frac{t^2}{2}}dt\text{ we get:}\\
a)\ P\{Z\leq z^*\}=0.95\\
\Phi(z^*)=0.95/2=0.475\\
z^*\approx 1.96\\
b)\ P\{Z\leq z^*\}=0.2\\
\Phi(z^*)=0.2/2=0.1\\
z^*\approx 0.25\\
c)\ P\{Z\leq z^*\}=0.25\\
\Phi(z^*)=0.25/2=0.125\\
z^*\approx 0.32\\
d)\ P\{Z\geq z^*\}=0.9\\
P\{Z\leq z^*\}=0.9-0.5\\
P\{Z\leq z^*\}=0.4\\
\Phi(z^*)=0.4\\
z^*\approx 1.28\\
\text{The answer is } \approx -1.28.\\
e)\ P\{0\leq Z\leq z^*\}=0.41\\
\Phi(z^*)=0.41\\
z^*\approx 1.34\\
f)\ P\{-z^*\leq Z\leq z^*\}=0.88\\
\Phi(z^*)=0.44\\
z^*\approx 1.55"